Suppose you sell six August 2017 gold futures contracts on this day, at the last price of the day. Use Table 23.1 a. What wilSample Wall Street Journal Futures Price Quotations TABLE 23.1 Metal & Petroleum Futures Contract Оpen interest High hi lo SeAgriculture Futures Corn (CBT)-5,000 bu; cents per bu. March 369.50 367-50 369.50 372.75 1.25 500,405 Мay 380.25 346,739 377.Interest Rate Futures Treasury Bonds (CBT)-$100,0o0; pts 32nds of 100 % Мarch 153-160 151-310 152-060 622,929 153-170 1-030 JIndex Futures Mini DJ Industrial Average (CBT)-$5 x index Мarch 20158 19982 20136 126,519 20010 134 June 136 893 19921 20093
Answer #1

1.
Loss of 3960.00
=6*100*(-1249.7+1243.1)=-3960.00

2.
Profit of 4320.00
=6*100*(-1235.9+1243.1)=4320.00